Farmer Sara has 14 more chickens than sheep. There is a total of 94 legs between the chickens and sheep. How many sheep does farmer Sara have?
Answer by Boreal(15235)

Sheep=S
chickens=S+14=C
legs are 4S+2C=94
but this is 4S+2(S+14)=4S+2S+28=94
6S=66
S=11 sheep or 44 legs ANSWER
C=25 chickens or 50 legs